A driveway is covered with concrete with an excess of 200 sq. ft. A walkway is to be placed around a flower garden 20 * 15 ft. Determine walkway width.
20 * 15 = 300 sq. ft.
300 / 200 = 1.5 ft.
Where am I correct? Thanks.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

IDK what that means.
=================
"Where am I correct?" makes no sense, either.
You were correct that 20*15 = 300
You were correct that 300/200 = 1.5
=====================
If you show a solution, you should ask, "Where am I wrong?"

Answer by ikleyn(52866)

.
A driveway is covered with concrete with an excess of 200 sq. ft. A walkway is to be placed around a flower garden 20 * 15 ft. Determine walkway width.
20 * 15 = 300 sq. ft.
300 / 200 = 1.5 ft.
Where am I correct? Thanks.
~~~~~~~~~~~~~~~~~~~~~~~~~~

1. "A driveway is covered with concrete with an excess of 200 sq. ft."

     This phrase doesn't make sense NEITHER from the Math point of view, NOR from the point of view of its content.


2. Two phrases in the condition are not relevant.
     The first is not relevant to the second.
     The second is not relevant to the first.


3. It seems to me that these phrases are mixed from two different problems.
     If any.


4. And where you ask "Where am I correct?",
     the answer is "NOTHERE".